Iron reacts with hydrochloric acid to produce iron (II) chloride and hydrogen gas:


Fe(s) + HCl(aq) -----> FeCl2 + H2 (g)


The H2 gas from the reaction of 300.0 g of iron with 300.0 g hydrochloric acid is collected in a 2000.mL flask at 20.00°C. What is the pressure of H2 gas in this flask in bars? (1 atm = 1.01325 bars)


1
Expert's answer
2021-12-07T02:06:02-0500

Balanced chemical equation:

Fe(s) + 2HCl(aq) → FeCl2(aq) + H2(g)

According to the equation above: n(Fe) = n(H2)


The molar mass of Fe is 55.845 g/mol.

Therefore,

(2.2 g Fe) × (1 mol Fe / 55.845 g Fe) = 0.0394 mol Fe


n(H2) = n(Fe) = 0.0394 mol


The Ideal Gas equation can be used.

The Ideal Gas equation can be expressed as: PV = nRT


P = unknown

V = 10.0 L

n = 0.0394 mol

R = 0.0821 atm L mol-1 K-1

T = 25˚C + 273 = 298 K


Thus,

P = nRT / V

P = (0.0394 mol × 0.0821 atm L mol-1 K-1 × 298 K) / (10.0 L) = 0.0964 atm

P = 0.0964 atm


Answer: 0.0964 atm is the pressure of H2 gas in this flask.